### v2

# VSI: Visual–Subtitle Integration for Keyframe Selection to Enhance Long Video Understanding

Abstract
Long video understanding presents a significant challenge to multimodal large language models (MLLMs) primarily due to the immense data scale. A critical and widely adopted strategy for making this task computationally tractable is keyframe retrieval, which seeks to identify a sparse set of video frames that are most salient to a given textual query. However, the efficacy of this approach is hindered by weak multimodal alignment between textual queries and visual content and fails to capture the complex temporal semantic information required for precise reasoning. To address this, we propose Visual-Subtitle Integration (VSI), a multimodal keyframe search method that integrates subtitles, timestamps, and scene boundaries into a unified multimodal search process. The proposed method captures the visual information of video frames as well as the complementary textual information through a dual-stream search mechanism by Video Search Stream as well as Subtitle Match Stream, respectively, and improves the keyframe search accuracy through the interaction of the two search streams. Experimental results show that VSI achieve 40.00% key frame localization accuracy on the text-relevant subset of LongVideoBench and 68.48% accuracy on downstream long Video-QA tasks, surpassing competitive baselines by 20.35% and 15.79%, respectively. Furthermore, on the LongVideoBench benchmark, VSI achieved state-of-the-art(SOTA) in medium-to-long video-QA tasks, demonstrating the robustness and generalizability of the proposed multimodal search strategy.
Introduction
Figure 1: Examples of Tasks Ineffectively Handled by Unimodal Methods (e.g.VSLS) but Effectively Processed by VSI. These tasks, sourced from LongVideoBench, include: Text-specified Object Attribute Localization (T2A), Text-specified Event Localization (T2E), Text-specified Object Localization (T2O), and Text-specified Object Tracking(TOS).
Breakthroughs in multimodal large language models (MLLMs) have driven the development of multimodal understanding. Among many MLLMs studies, video data processing has become an important direction. Compared with still images, videos contain more complex visual information, posing a serious challenge to MLLMs. Currently, the common way for MLLMs to process video information is to treat the video as continuous frames and encode the input frames into visual token sequences for processing by the MLLMs; however, when this framework is applied to long videos, the computational resource limitations result in the model only being able to process a very small portion of frames, and the quality of the sampled frames is closely related to the model’s effectiveness on the downstream tasks. Therefore, how to sample key frames to the instruction has become a critical problem to be solved.
Previous work has explored key-frame search algorithms that replace MLLMs with lightweight target detection models to process initial video frames and select keyframes for downstream Video-QA tasks. Tstar (Ye et al. 2025) uses YOLO (Cheng et al. 2024a) to identify key targets and increase the sampling weights of video frames containing them. Building on this, Visual Semantic-Logical Search (VSLS(Guo et al. 2025) mitigates the problem of missing temporal information by modeling relationships across video frames. While these approaches show progress in improving frame sampling, they fundamentally ignore the rich multimodal information embedded in videos, such as subtitles, scene text, or descriptive captions. They also overlook the semantic coherence between textual and visual cues. As a result, their performance stagnates on tasks requiring complex textual reasoning, limiting their applicability to text-related VideoQA tasks. To address this limitation, we propose Visual-Subtitle Integration (VSI), a multimodal keyframe search method. We design a dual-stream search path consisting of a Video Search stream and a Subtitle Matching stream, improving upon previous approaches. The Subtitle Matching stream uses textual similarity to align the question with subtitle content and localize relevant keyframes more accurately. This information is then fused with the Video Search stream to improve keyframe search accuracy.
VSI offers three key advantages: 1) Due to the complementary nature of textual and visual information, the framework can locate keyframes more efficiently and accurately, and achieve high-quality frame sampling; 2) It extends the keyframe search algorithm from unimodal to multimodal, making full use of visual as well as textual information, enabling the framework to handle text-related Video-QA tasks, which are not handled by unimodal search methods, and perform equally well on other types of tasks; 3) As a plug-and-play module, it does not require additional training, so it can be seamlessly integrated into existing video processing flows and flexibly adapted to different requirements by adjusting the object detection model and text encoding model.
We comprehensively evaluated VSI on several public benchmarks, including LongVideoBench, a large-scale benchmark for long video understanding, and Video-MME, a widely used dataset for multimodal video question answering. Experimental results demonstrate that for LongVideoBench-text-related perception subsets, VSI significantly improve keyframe search accuracy from 19.65% to 40.00%, representing a 20.35% absolute gain over the baseline. Furthermore, our approach boosts GPT-4o (Hurst et al. 2024) answer accuracy by 17.24% on medium-length video QA tasks and by 22.24% on long video-QA tasks. These substantial gains highlight the effectiveness of multimodal fusion in guiding frame selection and underscore the benefits of our lightweight optimization strategy.
Our principal contributions are threefold:
  • We extended the keyframe search algorithm from single-image modality search to multi-modality dual-stream search, enabling the algorithm to handle text-related problems that single-modality algorithms cannot handle, and it also performs excellently in other types of tasks.
  • We propose a plug-and-play method that requires no additional training and is both lightweight and flexible, with search models and encoding models that can be flexibly replaced according to different needs.
  • We conducted extensive empirical evaluations across multiple datasets. Experiments showed that with only four frames sampled, we achieved a keyframe search accuracy of 40% in text-related tasks and an accuracy rate of 68.48% in downstream Long Video-QA tasks.
Related work
Figure 2: Framework for Visual-Subtitle Integration. The dual-stream architecture comprises: (a) Video Search pathway leveraging YOLO-World to identify query-relevant objects; (b) Subtitle Match pathway employing contrastive embeddings to retrieve subtitle-matching segments; (c) Confidence scores from both modalities are fused to update frame-wise relevance probabilities through spline interpolation. After the iteration, high-confidence frames were subsequently propagated to downstream QA tasks. The figure shows a complete real example from keyframe search to the completion of a Video-QA task.
Challenges in Long-Form Video Understanding. Long-form video analysis (Xun et al. 2025) poses greater and more persistent challenges than short-video or image tasks (Yan et al. 2025), primarily due to inherently complex temporal dynamics and substantial content redundancy (Qian et al. 2024; Zeng et al. 2024; Yu et al. 2019; Dang et al. 2024). The large number of frames introduces significant computational burdens, substantially increasing memory and latency, making exhaustive frame-wise processing infeasible. Moreover, key semantic events are often temporally sparse and discontinuous, requiring specialized architectures capable of capturing subtle, long-range dependencies (Ranasinghe et al. 2025; Shi et al. 2024; Chen et al. 2024b; Weng et al. 2024). The visual complexity of such videos also introduces noisy, ambiguous, or irrelevant signals, which calls for more effective content distillation mechanisms (Zhang, Li, and Bing 2023; Cheng et al. 2024b; Xu et al. 2023; Ye et al. 2025; He et al. 2022).
Video-Language Modeling Strategies. Recent work using VLMs has primarily focused on three complementary directions: 1) content-aware sampling for more efficient computation (Buch et al. 2025; Yu et al. 2025; Hu et al. 2025), 2) hierarchical reasoning for complex query understanding (Wang et al. 2024a; Min et al. 2024; Liao et al. 2024b), and 3) information compression to better mitigate context window limitations (Tao et al. 2025; Liu et al. 2025; Choudhury et al. 2024; Chen et al. 2025). Representative strategies include segmentation-retrieval pipelines that divide long videos into clips and employ either learned or heuristic mechanisms to reliably select relevant segments (Pan et al. 2023; Choudhury et al. 2023, 2025), and fine-grained frame-level token compression to significantly reduce computational costs (Li, Wang, and Jia 2024; Chen et al. 2024a; Song et al. 2024), or adopt memory propagation techniques to better maintain temporal continuity (Qian et al. 2024; Wu et al. 2022; Liu et al. 2024). Recent advances also explore planning systems that decompose complex tasks into sequential perception steps, thereby enabling dynamic and adaptive frame-level information acquisition (Wang et al. 2024a; Liao et al. 2024a).
Method
Our proposed framework addresses the limitations of unimodal video search by introducing a Visual-Subtitle Integration (VSI) mechanism that effectively integrates visual-temporal search and textual semantic search for multimodal keyframe retrieval. Specifically, we construct two parallel search streams: (1) a video search stream that identifies salient frames based on spatiotemporal patterns, and (2) a subtitle matching stream that locates semantically relevant segments through contrastive sentence embeddings derived from subtitles. After each iteration, the text score produced by the subtitle matching stream is fused with the object score from the video search stream. The resulting fused score is then used to update the sampling probability distribution of the object detection model, allowing it to focus on semantically informative regions and reduce the average number of iterations. After the final iteration, the fused score distribution is used to select the top-$k$ keyframes as the final output.
Task Formulation
Given a video sequence $V=\{v_{i}\}_{i=1}^{T}$ consisting of $T$ frames and a textual query $Q$, the temporal retrieval task aims to identify the most concise keyframe subset $S=\{v_{k_{j}}\}_{j=1}^{K}\subset V$ that satisfies two criteria: 1) Answer Preservation: The extracted subset $S\subset V$ must guarantee answer equivalence, i.e., $F(S,Q)\sim F(V,Q)$, where $F(\cdot)$ denotes the video question-answering model; 2) Minimality: $S$ should form an irreducible subset that preserves the correctness of the answer—any frame removal should alter the model’s response—while maintaining computational efficiency.
Multi-modal Temporal-Semantic Fusion Search
Frame Sampling.
To expedite the search procedure, we employ a selective sampling approach rather than processing every video frame. Let $N_{\text{video}}$ represent the total frame count, and $D$ be an initial uniform sampling distribution across all frames. The frame selection is performed as:
$I_{\text{sel}}=\mathcal{R}(D\circ N_{\text{video}},n_{\text{sample}}),$ (1)
where $\mathcal{R}(\cdot,n_{\text{sample}})$ randomly picks $n_{\text{sample}}$ frames following the weighted distribution $D\circ N_{\text{video}}$. For computational efficiency, we arrange the selected frames into a $m\times m$ matrix format, which necessitates:
$n_{\text{sample}}=m^{2},\quad m\in\mathbb{N}^{+},\ \text{and}\ n_{\text{sample}}\ll N_{\text{video}}.$ (2)
This grid organization structure enables efficient batch processing. Although $D$ is initially uniformly distributed, it can be refined through iteration to prioritize more relevant frames in subsequent sampling processes.
Object Detection and Scoring.
To provide initial visual grounding, VSI prompts the VLM to scan a set of evenly spaced frames and identify relevant objects from two semantic categories: 1) Target Objects $T_{\text{obj}}$: key visual entities directly related to answering the query; 2) Cue Objects $C_{\text{obj}}$: contextual entities that offer indirect clues, such as common spatial configurations or environmental hints. We formalize this as:
$T_{\text{obj}},\,C_{\text{obj}}:=\text{VLMs}(V,Q,K),\quad\mathcal{T}:=T_{\text{obj}}\cup C_{\text{obj}}.$ (3)
During each sampling iteration, we apply an efficient detector (e.g., YOLO-World (Cheng et al. 2024a)) to the selected $n_{\text{sample}}$ frames. Each frame is evaluated by checking the overlap between detected objects and the predefined target set $\mathcal{T}$. Let $D_{t}$ denote the set of objects detected in frame $t$, $s_{o}$ the detection confidence of object $o\in D_{t}$, and $\alpha_{o}$ the task-specific importance weight for object $o$. The object-based frame score is defined as:
$S_{\text{obj}}(t)=\max_{o\in D_{t}\cap\mathcal{T}}\left(s_{o}\cdot\alpha_{o}\right).$ (4)
If $S_{\text{obj}}(t)$ exceeds a threshold $\tau$, the corresponding frame is added to a priority queue for downstream processing. This mechanism ensures that visually salient and semantically aligned frames are prioritized for further analysis.
Textual Semantic Search.
Algorithm 1 VSI $(V,Q,K,\tau)$
1:Video $V=\{v_{t}\}_{t=1}^{N_{v}}$, Query $Q$,
2:Keyframes $S=\{v_{k_{j}}\}_{j=1}^{K}$
[…]
5:while $\beta>0$ and $\neg\textsc{AllFound}(\mathcal{T})$ do
6:  $I_{s}\leftarrow\textsc{Sampling}(D,\lfloor\sqrt{\beta}\rfloor^{2})$ // Eq.(1)-(2)
7:  $S_{o}\leftarrow\textsc{ObjectScoring}(I_{s},\tau)$ // Eq.(3)-(4)
8:  $S_{t}\leftarrow\textsc{TextSearch}(Q,V_{sub})$ // Eq.(6)-(11)
9:  $S_{f}\leftarrow\textsc{ScoreFusion}(S_{o},S_{t},Q)$ // Eq.(12)-(13)
10:  $\textsc{UpdateDistribution}(D,S_{f})$ // Eq.(14)-(18)
11:  $\beta\leftarrow\beta-|I_{s}|$
12:end while
13:return $\textsc{TopK}(\mathcal{S}_{f},K)$ // Final keyframe selection
Given an input tuple comprising subtitle text segments $S=\left\{s_{1},s_{2},\ldots,s_{n}\right\}$ with corresponding timestamps $T=$ $\left\{t_{1},t_{2},\ldots,t_{n}\right\}$, and a textual query $Q$, our semantic search pipeline operates through the following computational steps: 1) Dual-Text Encoding: A pre-trained transformer encoder $E_{\text{text}}$ (specifically, all-mpnet-base-v2111https://huggingface.co/sentence-transformers/all-mpnet-base-v2) is used to encode both the textual query and subtitle segments into a shared $d$-dimensional latent space:
$\mathbf{q}=E_{\text{text}}(Q)\in\mathbb{R}^{d},$ (5)
$\mathbf{s}_{i}=E_{\text{text}}(s_{i})\in\mathbb{R}^{d},\quad\forall i\in[1,n],$ (6)
where $Q$ is the input query, $s_{i}$ denotes the $i$-th subtitle segment, and $d$ is the embedding dimension determined by the text encoder architecture; and 2) Similarity Computation: We calculate the cosine similarity matrix $\mathbf{C}\in\mathbb{R}^{n}$ between the query embedding and all subtitle embeddings, each element $\mathbf{C}_{i}$ is formulated as:
$\mathbf{C}_{i}=\frac{\mathbf{q}\cdot\mathbf{s}_{i}}{\|\mathbf{q}\|\left\|\mathbf{s}_{i}\right\|},\quad\forall i\in[1,n].$ (7)
Soft Threshold Enhancement: We apply soft threshold enhancement to the similarity scores $C_{i}$, using an amplification factor $\alpha$ and a threshold $\theta$:
$B_{i}=\min\left(C_{i}+\alpha\cdot\max(C_{i}-\theta,0),\ 1.0\right),$ (8)
where $B_{i}$ denotes the enhanced similarity score after soft thresholding, $\theta=0.5$ is the similarity threshold, $\alpha=2$ is the amplification factor, and $\min(\cdot,1.0)$ ensures that the score does not exceed 1.0.
Dynamic Gaussian Propagation: For each segment $i$ with $B_{i}>\beta$ ($\beta=0.2$), we propagate its score using a Gaussian kernel centered at the segment midpoint $c_{i}=(b_{i}+e_{i})/2$:
$w_{i}(t)=B_{i}\cdot\exp\left(-\frac{(t-c_{i})^{2}}{2\sigma_{i}^{2}}\right),$ (9)
where $b_{i}$ and $e_{i}$ are the begin and end timestamps of subtitle segment $i$, respectively. The standard deviation $\sigma_{i}$ controls the spread of the kernel and is defined as $\sigma_{i}=(e_{i}-b_{i}+2W)/{4}$, where $W=2$ is the temporal extension radius, and $(e_{i}-b_{i})$ is the original duration of the segment. The variable $t$ denotes a continuous time point within the video timeline.
Max-Confidence Aggregation: We compute the text confidence score $S_{\text{text}}(t)$ for the frame at time $t$ as the maximum propagated score among all relevant segments:
$S_{\text{text}}(t)=\max_{i:B_{i}>\beta,\,t\in[b_{i}-W,\ e_{i}+W]}w_{i}(t),$ (10)
with the convention that $\max(\emptyset)=0$, i.e., the score defaults to zero if no segment satisfies the condition. This operation aggregates the strongest textual signal across temporally adjacent segments.
Score Fusion Module
We fuse the object confidence score $S_{\text{obj}}(t)\in[0,1]$ and the text confidence score $S_{\text{text}}(t)\in[0,1]$ using an adaptive weighted scheme. Specifically, the fused score is computed as
$S_{\text{fused}}(t)=\alpha\cdot\mathcal{N}\left(S_{\text{text}}(t)\right)+\beta\cdot\mathcal{N}\left(S_{\text{obj}}(t)\right),$ (11)
where $\alpha+\beta=1$ are the modality weighting coefficients, and $\mathcal{N}(x)=(x-\mu_{x})/(\sigma_{x}+\epsilon)$ denotes the Z-score normalization, with $\mu_{x}$ and $\sigma_{x}$ being the temporal mean and standard deviation of score stream $x$. A small constant $\epsilon=1\times 10^{-6}$ ensures numerical stability.
Score Distribution Update
After each sampling iteration, we update the global score distribution $\{\mathcal{S}_{f}\}$ over all frames $f\in\{1,\cdots,N_{\text{video}}\}$. When a frame $f$ is selected for detection, its score is directly assigned:
$\mathcal{S}_{f}\leftarrow\operatorname{ConfidenceScore}(f),\quad\mathcal{V}\leftarrow\mathcal{V}\cup\{f\}.$ (12)
We then reconstruct scores for unvisited frames using spline interpolation over visited frames:
$\hat{\mathcal{S}}(f)=\operatorname{Spline}\left(\left\{(f_{i},\mathcal{S}_{f_{i}})\mid f_{i}\in\mathcal{V}\right\}\right)(f).$ (13)
To avoid degenerate probabilities, we apply a lower-bound correction:
$\tilde{\mathcal{S}}(f)=\max\left(N_{\text{video}}^{-1},\hat{\mathcal{S}}(f)\right).$ (14)
Finally, the frame selection distribution is obtained via a normalized sigmoid transformation:
$P_{f}=\frac{\sigma\left(\tilde{\mathcal{S}}(f)\right)}{\sum_{f^{\prime}=1}^{N_{\text{video}}}\sigma\left(\tilde{\mathcal{S}}(f^{\prime})\right)}.$ (15)
This process is repeated until the computational budget $\beta$ is reached or all targets are detected. The algorithm returns the top-$k$ frames ranked by $\mathcal{S}_{f}$.
In summary, the VSI framework integrates the complementary advantages of visual and textual modalities to provide a novel solution for multimodal video search algorithms.
Exeriments
Method Searching Modality Keyframe Source Text Weight Frame Acc $\uparrow$ LongVideoBench Medium Long GPT4o (Wu et al. 2024) NA Uniform Sampling N/A 4 N/A 46.55 46.24 GPT4o+Sub (Wu et al. 2024) NA Uniform Sampling N/A 4 N/A 44.83 45.16 GPT4o+Tstar (Ye et al. 2025) Unimodal Keyframe Searching N/A 4 19.65 53.45 53.76 GPT4o+VSLS (Guo et al. 2025) Unimodal Keyframe Searching N/A 4 18.50 50.00 52.69 GPT4o+VSI (ours) Multimodal Keyframe Searching 0.3 4 14.47 45.61 56.99 GPT4o+VSI (ours) Multimodal Keyframe Searching 0.7 4 18.99 55.17 54.95 GPT4o+VSI (ours) Multimodal Keyframe Searching 1.0 4 40.00 63.79 68.48 LLaVa-Video-7B (Wu et al. 2024) Unimodal Uniform Sampling N/A 4 N/A 34.48 26.88 LLaVa-Video-7B+VSI (ours) Multimodal Keyframe Searching 1.0 4 40.00 32.76 39.13 GPT4o (Wu et al. 2024) N/A Uniform Sampling N/A 8 N/A 48.28 53.76 GPT4o+Sub (Wu et al. 2024) N/A Uniform Sampling N/A 8 N/A 46.55 58.06 GPT4o+Tstar (Ye et al. 2025) Unimodal Keyframe Searching N/A 8 29.48 48.28 58.06 GPT4o+VSLS (Guo et al. 2025) Unimodal Keyframe Searching N/A 8 27.17 48.28 55.91 GPT4o+VSI (ours) Multimodal Keyframe Searching 0.3 8 20.75 45.61 56.99 GPT4o+VSI (ours) Multimodal Keyframe Searching 0.7 8 28.48 56.90 60.44 GPT4o+VSI (ours) Multimodal Keyframe Searching 1.0 8 45.00 62.07 69.57 LLaVa-Video-7B (Wu et al. 2024) Unimodal Uniform Sampling N/A 8 N/A 31.03 35.48 LLaVa-Video-7B+VSI (ours) Multimodal Keyframe Searching 1.0 8 45.00 36.21 36.96
Table 1: Keyframe Search Accuracy and Downstream QA Accuracy on LongVideoBench-Text-Related Perception Subsets. Where Searching Modality indicates whether the search algorithm uses unimodal or multimodal information, and Keyframe Source indicates whether the frames used for VideoQA come from the keyframe search algorithm or uniform sampling. Text Weight represents the weight of Text Score in performing score fusion.
Benchmark Datasets
Figure 3: Comparative Keyframe Search Accuracy between VSI and the unimodal baselines on LongVideoBench. Where TRP-kFrames denotes the case of keyframe search accuracy on Text-Related Perception Subset based on taking top-k frames, and Full denotes the full dataset.
We comprehensively evaluated the proposed VSI framework on two pivotal benchmarks: LongVideoBench (Wu et al. 2024) for assessing long-context video-language understanding capabilities, and Video-MME (Fu et al. 2024) as the first integrated benchmark for multimodal video analysis focused on QA evaluation. Leveraging ground-truth key frame annotations in LongVideoBench enables rigorous validation of key frame selection accuracy, while both datasets collectively support quantitative analysis of VideoQA performance, with detailed dataset statistics provided in the Appendix.
Evaluation of Subtitle Match Performance
To substantiate our method’s efficacy in leveraging textually relevant information, we evaluated keyframe search accuracy on the text-related perception subset of LongVideoBench. Specifically, we define the Text-Related Perception Subset as encompassing three task types: Text-referred Object Attribute (T2A), Text-referred Event (T2E), and Text-referred Object (T2O). Given that key frame localization in these tasks exhibits strong dependence on subtitle information, this subset provides direct empirical validation of our method’s superiority over unimodal approaches. Detailed dataset specifications are provided in the appendix.
Our method achieves state-of-the-art(SOTA) keyframe search accuracy on this subset. Keyframe search accuracy is defined as having at least one frame within the top-k highest-scoring frames deviating no more than 200 frames from the Ground Truth keyframe. Specifically, using only the top-4 candidate frames, VSI attains 40.00% search accuracy, surpassing the second-best method (19.65%) by 20.35%. For downstream QA tasks, when feeding only the top-4 keyframes into GPT-4o, our approach achieves 63.79% accuracy on medium-length video QA tasks(outperforming the best unimodal method’s 53.45% by 10.34%) and 68.84% accuracy on long video QA tasks(surpassing the second-best unimodal method’s 53.76% by 15.08%). This demonstrates that our method delivers high prediction accuracy with minimal frames, confirming the effectiveness of our caption utilization strategy. To further demonstrate the generalizability of our method, we also conducted experiments using LLaVa-Video-7B as the base model. Under the condition of inputting the top-4 frames, we improved LLaVa’s accuracy on the Long Video-QA task from 26.88% to 39.13%.
Notably, the LongVideoBench benchmark includes a subtitle-based baseline method: this approach uniformly samples k video frames(e.g., 8) and feeds their corresponding subtitles as prompts to GPT-4o (denoted GPT-4o+Sub). This baseline reaches 58.06% accuracy on long video QA tasks, exceeding caption-free methods by 4.30% and verifying the necessity of subtitles. In contrast, VSI using top-8 frames achieve 69.67% accuracy on the same task, outperforming this subtitle baseline by 11.51%. This conclusively demonstrates the superiority of our subtitle utilization mechanism.
Evaluation of Downstream Video QA Performance
LongVideoBench Video-MME Model and Size Searching Modality Frame Video Length Model and Size Searching Modality Frame Video Length Long Medium Long Medium 900-3600s 180-600s 30-60min 4-15min GPT-4o (Wu et al. 2024) N/A 4 46.81 48.79 GPT-4o N/A 4 54.75 55.06 GPT-4o + Tstar Unimodal 4 45.39 50.49 GPT-4o + Tstar Unimodal 4 54.64 59.33 GPT-4o + VSLS Unimodal 4 47.52 51.47 GPT-4o + VSLS Unimodal 4 56.76 60.00 GPT-4o + VSI (ours) Multimodal 4 50.09 51.94 GPT-4o + VSI (ours) Multimodal 4 54.75 58.20 GPT-4o (Wu et al. 2024) N/A 8 46.99 50.49 GPT-4o N/A 8 54.53 61.24 GPT-4o + Tstar Unimodal 8 47.34 52.94 GPT-4o + Tstar Unimodal 8 54.97 59.66 GPT-4o + VSLS Unimodal 8 49.11 52.21 GPT-4o + VSLS Unimodal 8 55.64 61.12 GPT-4o + VSI (ours) Multimodal 8 49.20 56.07 GPT-4o + VSI (ours) Multimodal 8 55.75 59.53 PLLaVA-34B+Sub N/A 32 49.10 50.80 GPT-4V+Sub N/A 10 56.9 59.7 LLaVA-Next-Video-34B+Sub N/A 128 45.90 48.70 Claude-3.5-Sonnet+Sub N/A 20 54.70 60.10 GPT-4o-mini+Sub N/A 250 53.40 56.90 LLaVA-OneVision-78B+Sub N/A 32 60.00 62.20 Qwen2-VL-7B+Sub N/A 256 52.50 56.70 TimeMarker-8B+Sub N/A 128 51.90 60.70 GPT-4o(0513)+Sub N/A 256 61.60 66.70 VideoChat-Flash-7B+Sub N/A 512 63.30 63.80
Table 2: Downstream QA task evaluation results on LongVideoBench and Video-MME. Frame indicates the number of frames used in performing VideoQA tasks. Meanwhile, below the main table, we add the performance of the current common LLMs (including the model size and frames used) on VideoQA tasks as a reference for comparison. The data is from the official websites of two datasets (for consistency with our 2-decimal precision, single-decimal values are shown with a trailing zero).
To conclusively demonstrate the superiority of VSI, we conducted downstream VideoQA performance tests on both LongVideoBench and Video-MME datasets. As shown in Table 2, videos are grouped by length into Medium, and Long (the two lengths of video are defined differently in the two datasets, as indicated in Table 2.).
In terms of keyframe search accuracy, as shown in Fig  3, regardless of the premise of taking top-4 frames or top-8 frames, VSI achieves SOTA on the LongVideoBench full dataset, with 29.81% and 38.40% keyframe search accuracy, respectively, which is an improvement of 1.10% and 0.33% compared with the second-best unimodal methods. It shows that our multimodal search framework is not only adaptable to keyframe search for text-related perception tasks, but also effective in keyframe search for other types of tasks, which reflects the generalization of VSI.
In terms of the correct rate of downstream QA tasks, for the LongVideoBench dataset, VSI achieved SOTA performance in both the medium video QA task as well as the long video QA task. When inputting only top-4 frames to the GPT4o for QA, our methods achieve 51.94% and 50.09% correct rates in the medium video QA task and long video QA task, respectively, which are higher than the baseline method by 3.15% and 3.28%, and also improve by 0.47% and 2.57% compared to the second-best unimodal method, respectively. The excellent performance of VSI on the LongVideoBench dataset illustrates the correlation between the correct selection of keyframes and the correct answers of the downstream tasks on the one hand, and the robustness of the VSI framework on the other. For Video-MME, since this dataset does not have annotations for ground truth frames, we are currently unable to test keyframe search accuracy on this dataset. However, VSI also performed well in downstream video-QA tasks. Based on inputting the top-8 frames, our method achieves SOTA performance on the long Video-QA task. Additionally, when inputting the top-4 frames, we improve GPT-4o’s answer accuracy rate on the Medium Video-QA task from 55.06% to 58.20%.
Analysis
Method Training Required Searching Efficiency Frame Acc $\uparrow$ Video Search Text Encoding Iteration $\downarrow$ Tstar-Attention(Ye et al. 2025) Training-Based N/A N/A N/A N/A N/A Tstar-Detector(Ye et al. 2025) Training-Free YOLO-World-110M N/A 43 4 27.74 VSLS-Detector(Guo et al. 2025) Training-Free YOLO-World-110M N/A 49 4 28.71 VSI-Detector(ours) Training-Free YOLO-World-110M all-mpnet-base-v2 26 4 29.81 Tstar-Attention(Ye et al. 2025) Training-Based N/A N/A N/A N/A N/A Tstar-Detector(Ye et al. 2025) Training-Free YOLO-World-110M N/A 43 8 36.84 VSLS-Detector(Guo et al. 2025) Training-Free YOLO-World-110M N/A 49 8 38.07 VSI-Detector(ours) Training-Free YOLO-World-110M all-mpnet-base-v2 26 8 38.40
Table 3: Evaluation of computational performance metrics across the LongVideoBench benchmark. Training Required shows whether the method needs to be trained, Video Search indicates the model used by the method for video modality processing, and Text Encoding indicates the model used by the method for text modality processing. Iteration represents the average number of iterations needed for each keyframe search.
Computational Efficiency Metrics
Prior research (Fan et al. 2024; Park et al. 2024; Wang et al. 2024b, c; Wu and Xie 2023) has predominantly focused on enhancing domain-specific accuracy metrics, often overlooking the computational overhead of temporal search processes. To comprehensively assess this aspect, our evaluation scheme introduces two quantitative measures:
1) Computational Complexity measured in FLOPs:
$\text{FLOPs}=\sum_{l=1}^{L}(2\cdot C_{l}^{in}\cdot C_{l}^{out}\cdot K_{l}^{2}\cdot H_{l}\cdot W_{l}+C_{l}^{out}\cdot H_{l}\cdot W_{l}),$ (16)
where $L$ denotes total layers, $C_{l}$ represents channels, $K_{l}$ the kernel size, and $H_{l}\times W_{l}$ the spatial dimensions at layer $l$.
Time Complexity
The proposed multimodal framework extends the original approach by incorporating textual semantic analysis alongside visual processing. The enhanced method maintains the original two-stage architecture while introducing a parallel text processing pipeline. Our complexity analysis focuses on the integrated workflow:
Initial Phase: The system simultaneously processes visual content through VLMs and analyzes subtitle text using a sentence transformer (e.g., all-mpnet-base-v2). This dual processing generates:1)Visual semantic set $\mathcal{S}_{v}$ (target/cue objects and visual relations) ;2)Textual semantic set $\mathcal{S}_{t}$ (temporal anchors and textual cues)
Search Phase: The algorithm now operates with an augmented semantic set $\mathcal{S}=\mathcal{S}_{v}\cup\mathcal{S}_{t}$, implementing a hybrid scoring mechanism:
$h^{\prime}(f,\mathcal{S})=h_{v}(f,\mathcal{S}_{v})+\alpha_{t}h_{t}(f,\mathcal{S}_{t}),$ (17)
where $\alpha_{t}$ controls text-visual balance.
Complexity Analysis: The multimodal search retains the original $\mathcal{O}(|\mathcal{S}|k\log n)$ YOLO-World detection complexity, with three key additions: 1) Text Processing Overhead: Subtitle analysis introduces $\mathcal{O}(m)$ complexity for $m$ text segments, which is negligible compared to visual processing; 2) Search Convergence: Text-guided sampling empirically reduces the number of iterations by 38%; 3) Fusion Cost: Cross-modal scoring adds a constant $\mathcal{O}(1)$ overhead per frame.
Bottleneck Preservation: YOLO-World detection remains the computational bottleneck. Our implementation shows only 9.7% increased inference time versus the visual-only baseline, demonstrating efficient multimodal integration.
Evaluation of Search Framework Efficiency
As shown in Table 3, based on the complementarity of textual and visual information, the frame scores provided by the Subtitle Match stream in VSI can assist in updating the frame sampling probability in the Video Search stream. Therefore, the object detection model (YOLO-World-110M used in this paper) can find all objects more quickly. Therefore, on the LongVideoBench dataset, compared to single-modal methods, the VSI achieve higher keyframe search accuracy. While the next-best method requires an average of 49 iterations, our method only needs an average of 26 iterations to complete a search process, demonstrating the efficiency and effectiveness of our method.
In order to serve VSI lightweight and efficient design purpose, we have also designed in the model selection: 1) In the video search model, we chose YOLO-World-110M (Cheng et al. 2024a). This is a real-time open vocabulary target detection model developed by Tencent AI Lab. YOLO-World-110M realizes real-time target detection with open vocabulary, and can locate objects with zero samples through natural language descriptions, its lightweight architecture combined with the heavy parameterization convolution technique also meets our lightweight needs; 2) In text coding model we chose all-mpnet-base-v2 (Reimers and Gurevych 2020), which is a sentence embedding model based on the pre-trained model microsoft/mpnet-base (Song et al. 2020). It demonstrates excellent robustness in standard semantic search scenarios. Its optimized contrastive learning mechanism calculates cosine similarity between similar texts more accurately and is good at handling text matching between sentences.
Conclusion
This paper presented Visual-Subtitle Integration (VSI), a dual-stream keyframe retrieval framework that integrates subtitle-based semantic matching with visual object detection to guide frame sampling for long-video understanding. By aligning the query with both time-stamped captions and detected objects, VSI updates the sampling distribution toward frames that are more relevant for reasoning, improving localization while keeping computation practical. Experiments show that VSI achieves SOTA performance on LongVideoBench both in terms of keyframe search accuracy and downstream Video-QA tasks, and the method shows consistent performance across different backbone architectures. Being training-free and plug-and-play, VSI can be easily integrated into existing video-language pipelines with minimal modifications.