### v3

# VSI: Visual–Subtitle Integration for Keyframe Selection to Enhance Long Video Understanding

Abstract
Multimodal large language models (MLLMs) demonstrate exceptional performance in vision-language tasks, yet their processing of long videos is constrained by input context length and high computational costs. Sparse frame sampling thus becomes a necessary preprocessing step, with sampled frame quality directly impacting downstream performance. Existing keyframe search algorithms achieve a balance between efficiency and sampled frame quality but heavily rely on the visual modality alone. This makes them difficult to adapt to text-related tasks and often leads to retrieval results deviating from core semantic content. To address this, we propose the Visual-Subtitle Integration (VSI), a multimodal keyframe retrieval framework. It employs a dual-branch collaborative retrieval approach combining Video Search and Subtitle Match to fuse complementary visual and textual information for precise localization. Experiments on LongVideoBench and VideoMME demonstrate that VSI achieves state-of-the-art accuracy in keyframe retrieval while delivering breakthrough performance in text-related tasks and exhibiting strong generalization across other tasks.
1 Introduction
Figure 1: Comparison with Baseline. In both medium and long Video Settings of LongVideoBench and Video-MME datasets, the performance of VSI method consistently outperforms the three baseline models (GPT-4o, LLaVA-Video-7B-Qwen2, and Qwen2.5-VL-7B-Instruct) using uniform sampling strategy.
Multimodal Large Language Models (MLLMs) have demonstrated exceptional capabilities across a diverse range of vision-language tasks, including image captioning, Video Question Answering (VideoQA), and Optical Character Recognition (OCR). These models achieve such strong performance by encoding visual information into model-compatible input representations. Upon extension to the video domain, video-based MLLMs have attracted considerable research interest in recent years. However, due to the inherent complex spatio-temporal dynamics of video data, which is different from that of static images, the processing of video data faces great challenges.
Constrained by computational resources and inherent model capabilities, existing MLLMs typically employ sparse frame sampling when processing video. However, different sampling strategies can lead to significant variations in video content analysis results, ultimately impacting MLLM performance on downstream tasks. Therefore, the core challenge lies in determining the optimal frame sampling strategy to maximize MLLM video understanding accuracy while minimizing computational overhead. Most existing MLLMs (e.g., LLaVa-Video [36]) employ uniform frame sampling for video processing, which frequently omits query-relevant critical information. Current research focuses on exploring keyframe extraction methods that require no training. For instance, LongVU [longvu] utilizes pre-trained feature extractors (like DINOv2-1B) to identify frames exhibiting significant cross-frame differences; CoS [cos] employs LLaVA-1.5-13B to filter query-relevant frames as input for Video-MLLM, though this approach introduces substantial computational overhead. AKS [aks] performs adaptive keyframe selection by comprehensively evaluating the relevance of keyframes to prompts and their coverage within the video. Another class of keyframe search algorithms attempts to replace MLLM processing of initial video frames with lightweight object detection models to screen keyframes for downstream video question-answering tasks. Tstar [ye2025rethinking] represents this approach by identifying key objects and boosting the sampling weights of frames containing them. Building upon this foundation, Visual Semantic Logic Search(VSLS[vsls] explicitly models relationships between video frames to mitigate the issue of missing temporal information. The aforementioned frameworks provide effective solutions for minimizing computational overhead while maintaining the accuracy of MLLM-based video understanding.
However, current keyframe search algorithms still face two fundamental challenges: (1) Limited task types can be handled. In VideoQA tasks, existing methods exhibit significant limitations: they can only effectively improve performance on visually-strong sub-tasks, while showing minimal improvement for text-strong sub-tasks, making it difficult to meet performance demands in multimodal scenarios; (2) Failure to make full use of multimodal information. Existing approaches rely solely on visual monomodal retrieval, lacking targeted guidance from modalities like text. This leads to keyframes being overly focused on visually dense regions, shifting the emphasis away from true semantic content and consequently resulting in low search accuracy.
To address the aforementioned challenges, we propose Visual-Subtitle Integration (VSI). This framework employs a dual-branch collaborative retrieval mechanism comprising Video Search and Subtitle Match: the Video Search branch performs initial keyframe sampling based on visual feature extraction and object detection, while the Subtitle Match branch obtains complementary textual information through semantic similarity calculations. Outputs from both branches undergo dynamic fusion strategies to achieve deep cross-modal information interaction. The core advantages of this dual-branch collaborative design manifest in two aspects: First, through the targeted guidance of textual modality information, VSI effectively handles the VideoQA subtask dominated by textual relevance, overcoming the scenario limitations of existing unimodal approaches and significantly expanding the applicability of keyframe retrieval algorithms. Second, the deep fusion of cross-modal information enables the framework to precisely locate keyframes containing core semantics, avoiding retrieval bias caused by excessive focus on visual information.
We comprehensively evaluated VSI on several public benchmarks, including LongVideoBench [lvb], a large-scale benchmark for long video understanding, and Video-MME [videomme], a widely used dataset for multimodal video question answering. Experimental results demonstrate that VSI outperforms the uniform sampling baseline across all settings and achieves higher keyframe search accuracy than unimodal keyframe search methods on all task types within the LongVideoBench dataset. Notably, on text-related tasks within LongVideoBench, VSI significantly improved keyframe search accuracy from the baseline of 19.65% to 40.00% when using a 4-frame setting. Furthermore, when key frames sampled by VSI were fed into the GPT-4o [hurst2024gpt] model for Long-VideoQA tasks, the model’s answer accuracy improved by 22.24% compared to the baseline.
Our principal contributions are threefold:
  • We extended the keyframe search algorithm from single-image modality search to multi-modality dual-branch search, enabling the algorithm to handle text-related problems that single-modality algorithms cannot handle, and it also performs excellently in other types of tasks.
  • We propose a plug-and-play method that requires no additional training and is both lightweight and flexible, with search models and encoding models that can be flexibly replaced according to different needs.
  • We conducted extensive empirical evaluations across multiple datasets. Experimental results validated its high efficiency and effectiveness, with the VSI method achieving an average keyframe search accuracy of 73.89% while requiring sampling of only 3.2% of the video frames.
2 Method
Figure 2: Framework for Visual-Subtitle Integration. The dual-branch architecture comprises: (a) Video Search branch leveraging YOLO-World to identify query-relevant objects; (b) Subtitle Match branch employing contrastive embeddings to retrieve subtitle-matching segments; (c) Confidence scores from both modalities are fused to update frame-wise relevance probabilities through spline interpolation. After the iteration, high-confidence frames were subsequently propagated to downstream QA tasks. The figure shows a complete real example from keyframe search to the completion of a VideoQA task.
This paper proposes a Visual-Subtitle Integration (VSI) mechanism that integrates visual-temporal and textual semantic retrieval to overcome unimodal limitations. It employs a dual-branch collaborative mechanism: (1) The video search branch samples keyframes based on query-relevant objects via object detection and assigns them high weights; (2) The subtitle matching branch enhances corresponding frame weights through query-caption semantic similarity. After initializing frame weights, the results fuse scores from both branches to update sampling probabilities, guiding sampling toward semantically dense regions. The top-$k$ keyframes are selected based on the score distribution.
2.1 Task Formulation
Given a video sequence $V=\{v_{i}\}_{i=1}^{T}$ consisting of $T$ frames and a textual query $Q$, the keyframe retrieval task aims to identify the most concise keyframe subset $S=\{v_{k_{j}}\}_{j=1}^{K}\subset V$ that satisfies two criteria: 1) Answer Preservation: The extracted subset $S\subset V$ must guarantee answer equivalence, i.e., $F(S,Q)\sim F(V,Q)$, where $F(\cdot)$ denotes the video question-answering model; 2) Minimality: $S$ should form an irreducible subset that preserves the correctness of the answer, with any frame removal altering the model’s response, while maintaining efficiency.
2.2 Visual-Subtitle Integration
2.2.1 Frame Sampling.
To expedite the search procedure, we employ a selective sampling approach rather than processing every video frame. Let $N_{\text{video}}$ represent the total frame count, and $D$ be an initial uniform sampling distribution across all frames. The frame selection is performed as:
$I_{\text{sel}}=\mathcal{R}(D\circ N_{\text{video}},n_{\text{sample}}),$ (1)
where $\mathcal{R}(\cdot,n_{\text{sample}})$ randomly picks $n_{\text{sample}}$ frames following the weighted distribution $D\circ N_{\text{video}}$. To further leverage the detector’s detection capabilities, we arrange the selected frames into a $m\times m$ matrix format, which necessitates:
$n_{\text{sample}}=m^{2},\quad m\in\mathbb{N}^{+},\ \text{and}\ n_{\text{sample}}\ll N_{\text{video}}.$ (2)
During the initial iteration, uniformly spaced sampling is employed to ensure comprehensive coverage of the video, establishing a preliminary understanding of its content while avoiding excessive early focus on any single region. Subsequently, $D$ adjusts through multiple sampling rounds to concentrate on frames of greater interest within the video.
2.2.2 Object Detection and Scoring.
To provide initial visual grounding, VSI prompts the VLM to scan a set of evenly spaced frames and identify relevant objects from two semantic categories: 1) Target Objects $T_{\text{obj}}$: key visual entities directly related to answering the query; 2) Cue Objects $C_{\text{obj}}$: contextual entities that offer indirect clues, such as common spatial configurations or environmental hints. Cue objects provide contextual clues to identify target object-related potential scenarios and can pinpoint relevant regions even in the absence of detected target objects, thereby boosting search efficiency. The specific prompt is provided in the appendix. We formalize this as:
$T_{\text{obj}},\,C_{\text{obj}}:=\text{VLMs}(V,Q,K),\quad\mathcal{T}:=T_{\text{obj}}\cup C_{\text{obj}}.$ (3)
During each sampling iteration, we apply an efficient detector (e.g., YOLO-World [yolo]) to the selected $n_{\text{sample}}$ frames. Each frame is evaluated by checking the overlap between detected objects and the predefined target set $\mathcal{T}$. Let $D_{t}$ denote the set of objects detected in frame $t$, $s_{o}$ the detection confidence of object $o\in D_{t}$, and $\alpha_{o}$ the task-specific importance weight for object $o$. The object-based frame score is defined as:
$S_{\text{obj}}(t)=\max_{o\in D_{t}\cap\mathcal{T}}\left(s_{o}\cdot\alpha_{o}\right).$ (4)
If $S_{\text{obj}}(t)$ exceeds a threshold $\tau$, the corresponding frame is added to a priority queue for downstream processing. This mechanism ensures that visually salient and semantically aligned frames are prioritized for further analysis.
2.2.3 Subtitle Matching
Figure 3: Case Study. (a)Sampling probability distribution of the Video Search branch at different iteration counts; (b)Influence of the Subtitle Match branch on the sampling probability distribution of the Video Search branch; (c)Timestamp corresponding to the subtitle with the highest similarity in the Subtitle Match branch. The three figures collectively illustrate how the textual semantic information provided by the Subtitle Match branch dynamically influences the sampling strategy of the Video Search branch.
Algorithm 1 VSI $(V,Q,K,\tau)$
1:Video $V=\{v_{t}\}_{t=1}^{N_{v}}$, Query $Q$,
2:Keyframes $S=\{v_{k_{j}}\}_{j=1}^{K}$
[…]
5:while $\beta>0$ and $\neg\textsc{AllFound}(\mathcal{T})$ do
6:  $I_{s}\leftarrow\textsc{Sampling}(D,\lfloor\sqrt{\beta}\rfloor^{2})$
7:  $S_{o}\leftarrow\textsc{ObjectScoring}(I_{s},\tau)$
8:  $S_{t}\leftarrow\textsc{TextSearch}(Q,V_{sub})$
9:  $S_{f}\leftarrow\textsc{ScoreFusion}(S_{o},S_{t},Q)$
10:  $\textsc{UpdateDistribution}(D,S_{f})$
11:  $\beta\leftarrow\beta-|I_{s}|$
12:end while
13:return $\textsc{TopK}(\mathcal{S}_{f},K)$ // Final keyframe selection
Given an input tuple comprising subtitle text segments $S=\left\{s_{1},s_{2},\ldots,s_{n}\right\}$ with corresponding timestamps $T=$ $\left\{t_{1},t_{2},\ldots,t_{n}\right\}$, and a textual query $Q$, our semantic search pipeline operates through the following computational steps: 1) Dual-Text Encoding: A pre-trained transformer encoder $E_{\text{text}}$ (specifically, all-mpnet-base-v2111https://huggingface.co/sentence-transformers/all-mpnet-base-v2) is used to encode both the textual query and subtitle segments into a shared $d$-dimensional latent space:
$\mathbf{q}=E_{\text{text}}(Q)\in\mathbb{R}^{d},$ (5)
$\mathbf{s}_{i}=E_{\text{text}}(s_{i})\in\mathbb{R}^{d},\quad\forall i\in[1,n],$ (6)
where $Q$ is the input query, $s_{i}$ denotes the $i$-th subtitle segment, and $d$ is the embedding dimension determined by the text encoder architecture; and 2) Similarity Computation: We calculate the cosine similarity matrix $\mathbf{C}\in\mathbb{R}^{n}$ between the query embedding and all subtitle embeddings, each element $\mathbf{C}_{i}$ is formulated as:
$\mathbf{C}_{i}=\frac{\mathbf{q}\cdot\mathbf{s}_{i}}{\|\mathbf{q}\|\left\|\mathbf{s}_{i}\right\|},\quad\forall i\in[1,n].$ (7)
Soft Threshold Enhancement: We apply soft threshold enhancement to the similarity scores $C_{i}$, using an amplification factor $\alpha$ and a threshold $\theta$:
$B_{i}=\min\left(C_{i}+\alpha\cdot\max(C_{i}-\theta,0),\ 1.0\right),$ (8)
where $B_{i}$ denotes the enhanced similarity score after soft thresholding, $\theta=0.5$ is the similarity threshold, $\alpha=2$ is the amplification factor, and $\min(\cdot,1.0)$ ensures that the score does not exceed 1.0.
Dynamic Gaussian Propagation: For each segment $i$ with $B_{i}>\beta$ ($\beta=0.2$), we propagate its score using a Gaussian kernel centered at the segment midpoint $c_{i}=(b_{i}+e_{i})/2$:
$w_{i}(t)=B_{i}\cdot\exp\left(-\frac{(t-c_{i})^{2}}{2\sigma_{i}^{2}}\right),$ (9)
where $b_{i}$ and $e_{i}$ are the begin and end timestamps of subtitle segment $i$, respectively. The standard deviation $\sigma_{i}$ controls the spread of the kernel and is defined as $\sigma_{i}=(e_{i}-b_{i}+2W)/{4}$, where $W=2$ is the temporal extension radius, and $(e_{i}-b_{i})$ is the original duration of the segment. The variable $t$ denotes a continuous time point within the video timeline.
Max-Confidence Aggregation: We compute the text confidence score $S_{\text{text}}(t)$ for the frame at time $t$ as the maximum propagated score among all relevant segments:
$S_{\text{text}}(t)=\max_{i:B_{i}>\beta,\,t\in[b_{i}-W,\ e_{i}+W]}w_{i}(t),$ (10)
With the convention that $\max(\emptyset)=0$, i.e., when no matching segments are found, the score defaults to 0. The aggregation operation under this agreement effectively integrates textual semantic information from adjacent segments across the temporal dimension, ultimately focusing on and extracting the strongest textual signals to achieve precise aggregation of key textual information at the temporal level.
2.2.4 Score Fusion Module
We fuse the object confidence score $S_{\text{obj}}(t)\in[0,1]$ and the text confidence score $S_{\text{text}}(t)\in[0,1]$ using an adaptive weighted scheme. The fused score is computed as:
$S_{\text{fused}}(t)=\alpha\cdot\mathcal{N}\left(S_{\text{text}}(t)\right)+\beta\cdot\mathcal{N}\left(S_{\text{obj}}(t)\right),$ (11)
where $\alpha+\beta=1$ are the modality weighting coefficients, and $\mathcal{N}(x)=(x-\mu_{x})/(\sigma_{x}+\epsilon)$ denotes the Z-score normalization, with $\mu_{x}$ and $\sigma_{x}$ being the temporal mean and standard deviation of score stream $x$. Introduce an infinitesimal constant $\epsilon=1\times 10^{-6}$ to ensure numerical stability.
2.2.5 Score Distribution Update
After each sampling iteration, we update the global score distribution $\{\mathcal{S}_{f}\}$ over all frames $f\in\{1,\cdots,N_{\text{video}}\}$. When a frame $f$ is selected for detection, its score is assigned:
$\mathcal{S}_{f}\leftarrow\operatorname{ConfidenceScore}(f),\quad\mathcal{V}\leftarrow\mathcal{V}\cup\{f\}.$ (12)
We then reconstruct scores for unvisited frames using spline interpolation over visited frames:
$\hat{\mathcal{S}}(f)=\operatorname{Spline}\left(\left\{(f_{i},\mathcal{S}_{f_{i}})\mid f_{i}\in\mathcal{V}\right\}\right)(f).$ (13)
This method can generate smooth probability curves that align with the continuity of video content. To avoid degenerate probabilities, we apply a lower-bound correction:
$\tilde{\mathcal{S}}(f)=\max\left(N_{\text{video}}^{-1},\hat{\mathcal{S}}(f)\right).$ (14)
Finally, the frame selection distribution is obtained via a normalized sigmoid transformation:
$P_{f}=\frac{\sigma\left(\tilde{\mathcal{S}}(f)\right)}{\sum_{f^{\prime}=1}^{N_{\text{video}}}\sigma\left(\tilde{\mathcal{S}}(f^{\prime})\right)}.$ (15)
This process is repeated until the computational budget $\beta$ is reached or target and cue objects are detected. The algorithm returns the top-$k$ frames ranked by $\mathcal{S}_{f}$.
In summary, the VSI framework provides a novel and effective solution for multimodal video keyframe retrieval by fully integrating the complementary characteristics of visual and textual modalities. To further validate the superiority of the proposed sampling strategy, this paper additionally conducts a comparative experiment, benchmarking the proposed adaptive iterative sampling strategy against the classic Epsilon-Greedy [e] and UCB [ucb] strategies. Experimental results demonstrate that compared to these two sampling strategies, the proposed adaptive iterative sampling strategy exhibits significant advantages: not only does it feature lower computational complexity and require fewer iterations per keyframe search, but it also achieves superior overall search efficiency. This outcome further validates the reliability and robustness of the proposed adaptive iterative sampling strategy. Detailed experimental data are in the Appendix.
3 Exeriments
Table 1: Downstream QA task evaluation results on LongVideoBench and Video-MME. Frame indicates the number of frames used in performing VideoQA tasks. Searching Modality indicates the primary information modalities utilized by keyframe search methods.
LongVideoBench Video-MME Model Searching Modality Frame Video Length Model Searching Modality Frame Video Length Long Medium Long Medium GPT-4O [hurst2024gpt] N/A 8 47.0 50.5 GPT-4O N/A 8 54.5 61.2 GPT-4O + Tstar [ye2025rethinking] Unimodal 8 47.3 52.9 GPT-4O + Tstar Unimodal 8 55.0 59.7 GPT-4O + VSLS [vsls] Unimodal 8 49.1 52.2 GPT-4O + VSLS Unimodal 8 55.6 61.1 GPT-4O + VSI (ours) Multimodal 8 49.2 56.1 GPT-4O + VSI (ours) Multimodal 8 55.8 59.5 LLaVA-Video-7B-Qwen2 [36] N/A 8 39.0 42.7 LLaVA-Video-7B-Qwen2 N/A 8 38.0 39.7 LLaVA-Video-7B-Qwen2+Tstar Unimodal 8 40.8 44.1 LLaVA-Video-7B-Qwen2+Tstar Unimodal 8 37.5 40.4 LLaVA-Video-7B-Qwen2+VSLS Unimodal 8 38.7 42.4 LLaVA-Video-7B-Qwen2+VSLS Unimodal 8 38.5 38.5 LLaVA-Video-7B-Qwen2+VSI (ours) Multimodal 8 40.6 44.1 LLaVA-Video-7B-Qwen2+VSI (ours) Multimodal 8 38.8 42.2 QWEN2.5-VL-7B-INSTRUCT [qwen] N/A 8 38.7 43.5 QWEN2.5-VL-7B-INSTRUCT N/A 8 38.0 47.3 QWEN2.5-VL-7B-INSTRUCT+Tstar Unimodal 8 39.2 43.9 QWEN2.5-VL-7B-INSTRUCT+Tstar Unimodal 8 40.3 50.0 QWEN2.5-VL-7B-INSTRUCT+VSLS Unimodal 8 42.4 43.9 QWEN2.5-VL-7B-INSTRUCT+VSLS Unimodal 8 43.2 49.6 QWEN2.5-VL-7B-INSTRUCT+VSI (ours) Multimodal 8 39.9 45.1 QWEN2.5-VL-7B-INSTRUCT+VSI (ours) Multimodal 8 40.8 48.5 GPT-4O N/A 32 48.2 51.9 GPT-4O N/A 32 55.2 61.0 GPT-4O + Tstar Unimodal 32 46.6 52.5 GPT-4O + Tstar Unimodal 32 55.2 61.6 GPT-4O + VSLS Unimodal 32 49.1 50.7 GPT-4O + VSLS Unimodal 32 57.5 61.9 GPT-4O + VSI (ours) Multimodal 32 51.2 53.9 GPT-4O + VSI (ours) Multimodal 32 55.3 61.7 LLaVA-Video-7B-Qwen2 N/A 32 39.4 45.3 LLaVA-Video-7B-Qwen2 N/A 32 35.9 36.5 LLaVA-Video-7B-Qwen2+Tstar Unimodal 32 39.2 43.9 LLaVA-Video-7B-Qwen2+Tstar Unimodal 32 35.8 39.6 LLaVA-Video-7B-Qwen2+VSLS Unimodal 32 42.4 43.9 LLaVA-Video-7B-Qwen2+VSLS Unimodal 32 36.9 39.0 LLaVA-Video-7B-Qwen2+VSI (ours) Multimodal 32 39.9 45.3 LLaVA-Video-7B-Qwen2+VSI (ours) Multimodal 32 38.2 42.2 QWEN2.5-VL-7B-INSTRUCT N/A 32 30.3 35.0 QWEN2.5-VL-7B-INSTRUCT N/A 32 37.5 39.9 QWEN2.5-VL-7B-INSTRUCT+Tstar Unimodal 32 38.7 42.9 QWEN2.5-VL-7B-INSTRUCT+Tstar Unimodal 32 34.9 45.6 QWEN2.5-VL-7B-INSTRUCT+VSLS Unimodal 32 38.1 43.1 QWEN2.5-VL-7B-INSTRUCT+VSLS Unimodal 32 37.9 50.0 QWEN2.5-VL-7B-INSTRUCT+VSI (ours) Multimodal 32 37.1 44.4 QWEN2.5-VL-7B-INSTRUCT+VSI (ours) Multimodal 32 39.3 45.8
3.1 Benchmark Datasets
We comprehensively evaluated the proposed VSI framework on two pivotal benchmarks: LongVideoBench [lvb] for assessing long-context video-language understanding capabilities, and Video-MME [videomme] as the first integrated benchmark for multimodal video analysis focused on QA evaluation. Leveraging ground-truth key frame annotations in LongVideoBench enables rigorous validation of key frame selection accuracy, while both datasets collectively support quantitative analysis of VideoQA performance, with detailed dataset statistics provided in the Appendix.
3.2 Search effectiveness and efficiency
For the quantitative evaluation of search performance, we adopt keyframe search accuracy as the core metric. Given that the LongVideoBench dataset provides authoritative ground truth annotations for key frames, we define this accuracy mathematically as follows:
Let the ground truth keyframe index of a video be $g$, and the set of top-k frames selected based on score distribution be $F=\{f_{1},f_{2},\ldots,f_{k}\}$ where $f_{i}$ denotes the index of the $i$-th frame in the top-k result. A search is deemed valid if there exists at least one frame in $F$ such that the frame index deviation from $g$ does not exceed 200 frames, i.e.,
$\exists f\in F\quad\text{s.t.}\quad|f-g|\leq 200$ (16)
Let $N_{\text{valid}}$ denote the number of videos with valid searches, and $N_{\text{total}}$ denote the total number of searched videos. The keyframe search accuracy is then defined as the proportion of validly searched videos relative to the total number of searched videos:
$\text{Accuracy}=\frac{N_{\text{valid}}}{N_{\text{total}}}$ (17)
This definition ensures that the retrieved frames still retain the core semantics of the ground truth (GT) keyframe. Specifically, these proximate frames adequately enable subsequent VLMs to comprehend critical video content, thus avoiding the loss of utility of retrieval results for downstream tasks caused by excessive frame deviation.
To conduct a systematic quantitative evaluation of search efficiency, this paper employs three core metrics: TFLOPs, Iteration, and Latency(sec). Specifically, TFLOPs (Trillions of Floating-Point Operations Per Second) quantifies the computational complexity of the search process; Iteration denotes the average number of iterations required to complete a single keyframe search task; and Latency represents the average time elapsed from the initiation to the completion of a single keyframe search task.
In this study, Tstar and VSLS—two approaches that also employ a keyframe search architecture—are selected as baseline models. As shown in Table 2, under the top-64 frame setting, the VSI method achieves state-of-the-art (SOTA) performance in keyframe search accuracy on the full LongVideoBench dataset, as well as on its image-only subset and text-only subset individually. Specifically, compared to the second-best performing methods, VSI demonstrates accuracy improvements of 3.66%, 1.22%, and 10.81% across these three evaluation scenarios, respectively. The improvement observed on the text-only subset aligns with intuitive expectations, indicating that the proposed method effectively leverages textual modality information to enhance search accuracy in this context. Notably, VSI also outperforms unimodal keyframe search methods on the image-only subset alone, suggesting that the subtitle matching stream provides complementary textual information that supports the video search stream, thereby significantly boosting keyframe search accuracy. This outcome underscores the robustness and adaptability of the VSI framework across diverse query scenarios. Furthermore, VSI achieves these performance gains without incurring significant increases in computational complexity, as measured by TFLOPs, average iteration count, or search latency, thus maintaining an efficient balance between retrieval effectiveness and efficiency.
3.3 Downstream VideoQA Performance
Table 2: Evaluation of computational performance metrics across the LongVideoBench benchmark. Training Required shows whether the method needs to be trained, Video Search indicates the model used by the method for video modality processing, and Text Encoding indicates the model used by the method for text modality processing.
Searching Efficiency Method Training Required Video Search Text Encoding TFLOPs Iteration Latency(sec) Frame Acc $\uparrow$ ALL T*-DETECTOR [ye2025rethinking] Training Free YOLO-World-110M N/A 31.7 24.76 28.96 64 67.58 VSLS-DETECTOR [vsls] Training Free YOLO-World-110M N/A 33.3 24.49 33.26 64 70.23 VSI-DETECTOR(ours) Training Free YOLO-World-110M ALL-MPNET-BASE-V2 36.8 26.34 31.71 64 73.89 IMAGE-ONLY T*-DETECTOR Training Free YOLO-World-110M N/A 31.7 23.90 25.4 64 70.34 VSLS-DETECTOR Training Free YOLO-World-110M N/A 33.3 24.04 28.13 64 68.62 VSI-DETECTOR(ours) Training Free YOLO-World-110M ALL-MPNET-BASE-V2 36.8 25.35 30.92 64 71.56 TEXT-ONLY T*-DETECTOR Training Free YOLO-World-110M N/A 31.7 26.80 27.99 64 66.36 VSLS-DETECTOR Training Free YOLO-World-110M N/A 33.3 26.39 28.77 64 66.36 VSI-DETECTOR(ours) Training Free YOLO-World-110M ALL-MPNET-BASE-V2 36.8 28.37 34.44 64 77.17
To conclusively demonstrate the superiority of VSI, we conducted downstream VideoQA performance tests on both LongVideoBench (LVB) and Video-MME (VMME) datasets. We focus on the performance on Medium(LVB:3-10min,VMME:4-15min) and Long(LVB:10-60min,VMME:30-60min) videos.
For the full-scale VideoQA task, we selected GPT-4o, LLaVA-Video-7B-Qwen2, and Qwen2.5-VL-7B-instruct as foundation models. As shown in Table 1, experiments were conducted with two input frame configurations: 8 frames and 32 frames. Experimental results demonstrate that for medium-length video tasks on LVB, VSI achieves state-of-the-art (SOTA) performance under both input frame configurations. Furthermore, across all experimental settings, VSI consistently exhibits significant performance improvements compared to baseline methods employing uniform sampling strategies, fully validating its effectiveness in medium-to-long video QA scenarios.
3.4 Visual Subtitle Integration Performance
To substantiate our method’s efficacy in leveraging textually relevant information, we evaluated keyframe search accuracy on the text-related perception subset of LongVideoBench. Specifically, we define the Text-Related Perception Subset as encompassing three task types: Text-referred Object Attribute (T2A), Text-referred Event (T2E), and Text-referred Object (T2O). Given that key frame localization in these tasks exhibits strong dependence on subtitle information, this subset provides direct empirical validation of our method’s superiority over unimodal approaches. Detailed dataset specifications are provided in the appendix. As shown in Table 3, VSI reaches 40% key frame search accuracy with only top-4 frames, a 20.35% gain over the second-best unimodal method. For downstream VideoQA’s medium and long video tasks, it achieves 63.79% and 68.48% accuracy, improving by 10.34% and 14.72% over the second-best method.
Notably, the LongVideoBench benchmark includes a subtitle-based baseline method: this approach uniformly samples k video frames(e.g., 8) and feeds their corresponding subtitles as prompts to GPT-4o (denoted GPT-4o+Sub). This baseline reaches 58.06% accuracy on long video QA tasks, exceeding caption-free methods by 4.30% and verifying the necessity of subtitles. In contrast, VSI using top-8 frames achieve 69.67% accuracy on the same task, outperforming this subtitle baseline by 11.51%.
Table 3: Keyframe Search Accuracy and Downstream QA Accuracy on LongVideoBench-Text-Related Perception Subsets.
Method Frame Acc $\uparrow$ LongVideoBench Medium Long GPT4o [lvb] 4 N/A 46.55 46.24 GPT4o+Sub [lvb] 4 N/A 44.83 45.16 GPT4o+Tstar [ye2025rethinking] 4 19.65 53.45 53.76 GPT4o+VSLS [vsls] 4 18.50 50.00 52.69 GPT4o+VSI (ours) 4 40.00 63.79 68.48 GPT4o [lvb] 8 N/A 48.28 53.76 GPT4o+Sub [lvb] 8 N/A 46.55 58.06 GPT4o+Tstar [ye2025rethinking] 8 29.48 48.28 58.06 GPT4o+VSLS [vsls] 8 27.17 48.28 55.91 GPT4o+VSI (ours) 8 45.00 62.07 69.57
4 Analysis
4.1 Video Search and Subtitle Match
To further quantify the independent contributions of the Video Search and Subtitle Match branches, we also tested and reported the performance of the standalone Video Search and Subtitle Match branches on the image-only and text-only subsets of the LongVideoBench dataset. As shown in Table 4, keyframe search accuracy results indicate that across all experimental settings on the image-only subset, the Video Search branch consistently achieves higher search accuracy than the Subtitle Match branch. Conversely, on the text-only subset, the Subtitle Match branch demonstrates significantly superior search performance compared to the Video Search branch. This branch also exhibits excellent adaptability in the downstream VideoQA task.
From the perspective of dual-branch interaction, during the keyframe search phase, the Video Search branch undertakes the core search task and serves as the primary module for completing keyframe retrieval. The Subtitle Match branch plays a supporting role, primarily injecting text-based semantic information that complements the image modality into the Video Search branch. Assisting this branch in dynamically updating its sampling probability distribution enables more precise targeting of video segments containing core semantic information. In the final keyframe selection phase, since keyframe selection relies on the combined score distribution output by both branches, the contribution weight of each branch to this distribution is, to a certain extent, determined by the relevance match between the current VideoQA query and either the image modality or the text modality.
As shown in Figure 3, the video timestamp corresponding to the subtitle with the highest semantic similarity matched by the Subtitle Match branch is 127.79s–127.80s. Subsequently, through the Score Fusion mechanism, this key textual semantic information is injected into the sampling probability distribution of the Video Search branch. This process enables the branch to focus more precisely on video frame regions containing core information relevant to the query during iterative searches, fully demonstrating the guiding role of text-modal information in visual search.
Table 4: Ablation Study of Video Search branch and Subtitle Match branch. TEXT-ONLY and IMAGE-ONLY refer to tasks within the Text-Referred category on LongVideoBench and all other task types. VideoQA uses GPT4o as the base model.
Method Frame Acc $\uparrow$ LongVideoBench Long Medium IMAGE-ONLY VSI-Subtitle Match 4 20.18 49.56 51.91 VSI-Video Search 4 23.55 48.09 52.67 VSI-Subtitle Match 8 27.68 50.15 54.58 VSI-Video Search 8 34.56 49.27 59.16 TEXT-ONLY VSI-Subtitle Match 4 40.00 68.48 60.34 VSI-Video Search 4 14.37 60.22 53.45 VSI-Subtitle Match 8 45.00 69.57 62.07 VSI-Video Search 8 20.00 58.06 55.17
4.2 Ablation Study of Text Weight
To further quantify the independent utility of the Subtitle Match Branch, we conduct ablation experiments targeting the core parameter Text Weight. This parameter’s primary function is to dynamically adjust the weighting proportion of the Subtitle Match Branch’s output scores during the score fusion stage of keyframe search. As shown in Table 5, experiments on the Text-Related Perception Subsets of the LongVideoBench dataset reveal a significant pattern: As Text Weight progressively increases, both keyframe search accuracy and answer correctness in the downstream VideoQA task show synchronous improvement. This dynamic outcome fully validates that when addressing highly text-related VideoQA problems, the Subtitle Match Branch provides crucial text-modal support. Its presence is indispensable, further confirming the rationality of this design.
Table 5: Keyframe Search Accuracy and Downstream QA Accuracy on LongVideoBench-Text-Related Perception Subsets.
Method Text Weight Frame Acc $\uparrow$ LongVideoBench Medium Long GPT4o+VSI 0.3 4 14.47 45.61 56.99 GPT4o+VSI 0.7 4 18.99 55.17 54.95 GPT4o+VSI 1 4 40.00 63.79 68.48 GPT4o+VSI 0.3 8 20.75 45.61 56.99 GPT4o+VSI 0.7 8 28.48 56.9 60.44 GPT4o+VSI 1 8 45.00 62.07 69.57
4.3 Time Complexity
The proposed multimodal framework extends the unimodal approach by incorporating textual semantic analysis alongside visual processing. The enhanced method maintains the original two-stage architecture while introducing a parallel text processing pipeline. The multimodal search retains the original $\mathcal{O}(|\mathcal{S}|k\log n)$ YOLO-World detection complexity, with two key additions: 1) Text Processing Overhead: Subtitle analysis introduces $\mathcal{O}(m)$ complexity for $m$ text segments, which is negligible compared to visual processing; 2) Fusion Cost: Cross-modal scoring adds a constant $\mathcal{O}(1)$ overhead per frame. YOLO-World detection remains the computational bottleneck. Our implementation shows only 10.5% increased inference time versus the visual-only baseline, demonstrating efficient multimodal integration.
5 Related work
Challenges in Long-Form Video Understanding. Long-form video analysis [xun2025rtv, liu2025javisgpt] poses greater and more persistent challenges than short-video or image tasks [survey], primarily due to inherently complex temporal dynamics and substantial content redundancy [qian2024streaming, zeng2024timesuite, yu2019activityqa, dang2024exploring]. The large number of frames introduces significant computational burdens, substantially increasing memory and latency, making exhaustive frame-wise processing infeasible. Moreover, key semantic events are often temporally sparse and discontinuous, requiring specialized architectures capable of capturing subtle, long-range dependencies [ranasinghe2025understanding, shi2024unlocking, chen2024rextime, weng2024longvlm]. The visual complexity of such videos also introduces noisy, ambiguous, or irrelevant signals, which calls for more effective content distillation mechanisms [84, cheng2024videollama2advancingspatialtemporal, xu2023retrieval, ye2025rethinking, hjx, fmri, d].
Video-Language Modeling Strategies. Recent work using VLMs has primarily focused on three complementary directions: 1) content-aware sampling for more efficient computation [buch2025flexible, yu2025framevoyager, hu2025mllm, tspo, moss], 2) hierarchical reasoning for complex query understanding [wang2024videoagent, min2024morevqa, liao2024align], and 3) information compression to better mitigate context window limitations [tao2025dycoke, liu2025hicom, choudhury2024rlt, chen2025longvila]. Representative strategies include segmentation-retrieval pipelines that divide long videos into clips and employ either learned or heuristic mechanisms to reliably select relevant segments [pan2023retrieving, choudhury2023zero, rohan2025video], and fine-grained frame-level token compression to significantly reduce computational costs [li2024llamavid, chen2024llavolta, song2024less], or adopt memory propagation techniques to better maintain temporal continuity [qian2024streaming, wu2022memvit, liu2024llava]. Recent advances also explore planning systems that decompose complex tasks into sequential perception steps, thereby enabling dynamic and adaptive frame-level information acquisition [wang2024videoagent, liao2024videoinsta].
6 Conclusion
This paper proposes Visual-Subtitle Integration (VSI), a multimodal keyframe search framework, whose core innovation lies in a dual-branch cross-modal collaborative retrieval mechanism combining Video Search and Subtitle Match. Through deep fusion of information from both branches, VSI achieves semantic-level precise localization of keyframes. On the LongVideoBench dataset, VSI demonstrates highly efficient retrieval performance: achieving 73.89% keyframe retrieval accuracy while sampling only 3.2% of video frames on average, enabling VLMs to outperform the uniform sampling baseline in downstream VideoQA performance. Moreover, VSI demonstrates adaptability and robustness in text-related tasks. VSI operates as a plug-and-play solution requiring no additional training or annotation, enabling seamless integration into existing video processing pipelines with outstanding engineering practicality.